OpenStudy (anonymous):

PLZ HELP Find 2 consecutive whole numbers that √74 lies between

OpenStudy (anonymous):

@rational help?

OpenStudy (rational):

first, try to find two perfect squares that 74 lies between

OpenStudy (rational):

whats the perfect square near to 74 that comes to your mind ?

OpenStudy (anonymous):

1,2 and 37

OpenStudy (anonymous):

@rational

OpenStudy (rational):

not exactly, a perfect square is number like : 4, 9, 16, 25, 36 etc... its a square of an integer

OpenStudy (rational):

the perfect squares near to 74 are 64 and 81 because 64 = 8^2 and 81 = 9^2

OpenStudy (rational):

64 < 74 < 81 8^2 < 74 < 9^2 8 < sqrt(74) < 9

OpenStudy (rational):

so sqrt(74 lies between the whole numbers 8 and 9